OpenStudy (anonymous):

HELPPPPP find the slope of a line perpendicular -x-3-y=0

OpenStudy (anonymous):

First I'd suggest changing the form of the equation into slope intercept form (y=mx+b). Then it'll be easier to work with.

OpenStudy (anonymous):

Do u know the property of two perpendicular lines?

OpenStudy (anonymous):

Put the equation in the form of y=mx+c

OpenStudy (anonymous):

I tried putting it into y = mx + b but I don't know if its right

OpenStudy (anonymous):

so a perpendicular line satisfies the condition that it's slope is the negative reciprocal of the initial line's slope. once you have the initial equation in y=mx+b then you can easily determine the slope of the initial line. then find the new slope by taking the negative reciprocal of the original slope.

OpenStudy (anonymous):

so do I plug in the 3 for b?

OpenStudy (anonymous):

0 = -x - 3?

OpenStudy (anonymous):

where did the 0 come from? you're close. you should add y to each side to get y= instead of 0 equals.

OpenStudy (anonymous):

y = 0

OpenStudy (anonymous):

Just isolate "y" and put every damn thing to the other side of the equation but remember y should be positive

OpenStudy (anonymous):

not quite.. so you start with the equation -x-3-y=0 and you want the equation to be in y= form. right now you have negative y on the left side, so the easiest way to get it into y= form is to add y to each side:  now its in y = form so whats the slope?